OpenStudy (ilovehorses):

5.Create your own equation written in standard form.

OpenStudy (ilovehorses):

@AnswerMyQuestions

OpenStudy (anonymous):

Hmm...

OpenStudy (ilovehorses):

hold on ill get the rest of the question

OpenStudy (ilovehorses):

Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (anonymous):

Sorry, I don't know that one. :(

OpenStudy (ilovehorses):

@supremesmack @CharleeAnne

OpenStudy (studygurl14):

Okay, just write any old equation for a line, using the form y = mx + b then pick any point, and determine if it is on the line by plugging it into your equation. :)

OpenStudy (anonymous):

idk

OpenStudy (ilovehorses):

I don't really understand

OpenStudy (ilovehorses):

@StudyGurl14 @Isaiah250 @CharleeAnne

OpenStudy (anonymous):

...

OpenStudy (studygurl14):

First step: Write an equation for a line such as... y = 2x + 4 This pick a point, like (1, 4) Plug it into equation (4) = 2(1) + 4 4 = 2 + 4 4 = 6 Not correct, so (1,4) isn't a point on the line See?

OpenStudy (studygurl14):

Do you know how to do this part: "Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts."

OpenStudy (ilovehorses):

not really

OpenStudy (anonymous):

byeeeeee

OpenStudy (studygurl14):

X and y intercepts: http://www.purplemath.com/modules/intrcept.htm

OpenStudy (ilovehorses):

ok ty for all the help im your fan and gave u a medal

OpenStudy (ilovehorses):

@StudyGurl14 could you come back and help me plz I need to know how do I find the correct point for the equation

OpenStudy (ilovehorses):

@emily_96 please help somebody

OpenStudy (studygurl14):

Sorry @ILOVEHORSES I have to go

OpenStudy (ilovehorses):

ok bye

OpenStudy (anonymous):

so what is your question? do you just want to make an equation in standard form?

OpenStudy (ilovehorses):

yes then answer these Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (anonymous):

well the standard form of a linear equation is Ax+By+C=0 so thats the form your answer should be in, since you have to make an equation yourself so that means you are free to choose any point you would like. so I would start off this question using the slope intercept form to make it easier This is in "y=mx+b" form; m = slope of your equation(the rate at which your equation is growing) b= vertical shift (how many units above or below the x-axis your equation is) so pick a number for the slope (m) pick a number for the vertical shift (b) first